### v3

# DifFace: Blind Face Restoration with Diffused Error Contraction

Abstract
While deep learning-based methods for blind face restoration have achieved unprecedented success, they still suffer from two major limitations. First, most of them deteriorate when facing complex degradations out of their training data. Second, these methods require multiple constraints, e.g., fidelity, perceptual, and adversarial losses, which require laborious hyper-parameter tuning to stabilize and balance their influences. In this work, we propose a novel method named DifFace that is capable of coping with unseen and complex degradations more gracefully without complicated loss designs. The key of our method is to establish a posterior distribution from the observed low-quality (LQ) image to its high-quality (HQ) counterpart. In particular, we design a transition distribution from the LQ image to the intermediate state of a pre-trained diffusion model and then gradually transmit from this intermediate state to the HQ target by recursively applying a pre-trained diffusion model. The transition distribution only relies on a restoration backbone that is trained with $L_{1}$ loss on some synthetic data, which favorably avoids the cumbersome training process in existing methods. Moreover, the transition distribution can contract the error of the restoration backbone and thus makes our method more robust to unknown degradations. Comprehensive experiments show that DifFace is superior to current state-of-the-art methods, especially in cases with severe degradations. Code and model are available at https://github.com/zsyOAOA/DifFace.
Index Terms:
Blind face restoration, face inpainting, diffusion model, error contraction.
I Introduction
Blind face restoration (BFR) aims at recovering a high-quality (HQ) image from its low-quality (LQ) counterpart, which usually suffers from complex degradations, such as noise, blurring, and downsampling. BFR is an extremely ill-posed inverse problem as multiple HQ solutions may exist for any given LQ image.
Approaches for BFR have been dominated by deep learning-based methods [1, 2, 3, 4]. The main idea of them is to learn a mapping, usually parameterized as a deep neural network, from the LQ images to the HQ ones based on a large amount of pre-collected LQ/HQ image pairs. In most cases, these image pairs are synthesized by assuming a degradation model that often deviates from the real one. Most existing methods are sensitive to such a deviation and thus suffer a dramatic performance drop when encountering mismatched degradations in real scenarios.
Figure 1: Overview of the proposed method. The solid lines denote the whole inference pipeline of our method. For ease of comparison, we also mark out the forward and reverse processes of the diffusion model by dotted lines.
Various constraints or priors have been designed to mitigate the influence of such a deviation and improve the restoration quality. The $L_{2}$ (or $L_{1}$) loss is commonly used to ensure fidelity, although these pixel-wise losses are known to favor the prediction of an average (or a median) over plausible solutions. Recent BFR methods also introduce the adversarial loss [5] and the perceptual loss [6, 7] to achieve more realistic results. Besides, some existing methods also exploit face-specific priors to further constrain the restored solution, e.g., face landmarks [8], facial components [9], and generative priors [10, 11, 1, 12, 13]. Considering so many constraints together makes the training unnecessarily complicated, often requiring laborious hyper-parameter tuning to make a trade-off among these constraints. Worse, the notorious instability of adversarial loss makes the training more challenging.
Figure 2: Comparative results of the proposed method to recent state-of-the-art approaches on the task of blind face restoration (top row) and face image inpainting (bottom row). Note that the masked area is highlighted using a purple color in the example of face inpainting.
In this work, we devise a novel BFR method DifFace, inspired by the recent success of diffusion model in image generation [14, 15]. Our method does not require multiple constraints in training, and yet is capable of dealing with unknown and complex degradations. Importantly, we leverage the rich image priors and strong generative capability of a pretrained diffusion model without retraining it on any manually assumed degradations.
To this end, we establish a posterior distribution $p(\bm{x}_{0}|\bm{y}_{0})$, aiming to infer the HQ image $\bm{x}_{0}$ conditioned on its LQ counterpart $\bm{y}_{0}$. Due to the complex degradations, solving for this posterior is non-trivial in blind restoration. Our solution to this problem, as depicted in Fig. 1, is to approximate this posterior distribution by a transition distribution $p(\bm{x}_{N}|\bm{y}_{0})$, where $\bm{x}_{N}$ is a diffused version of the desirable HQ image $\bm{x}_{0}$, followed with a reverse Markov chain that estimates $\bm{x}_{0}$ from $\bm{x}_{N}$. The transition distribution $p(\bm{x}_{N}|\bm{y}_{0})$ can be built up by introducing a neural network that is trained just using $L_{1}$ loss, and the reverse Markov chain is readily borrowed from a pretrained diffusion model containing abundant image priors.
The constructed transition distribution is appealing in that it is well motivated by an important observation in DDPM [16], where data is destroyed by re-scaling it with a factor of less than 1 and adding noise in the diffusion process. Bringing this notion to our context, the residual between $\bm{x}_{0}$ and $\bm{y}_{0}$ is also contracted by this factor after diffusion. Our framework uniquely leverages this property by inferring the intermediate diffused variable $\bm{x}_{N}$ (where $N<T$) from the LQ image $\bm{y}_{0}$, of which the residual to HQ image $\bm{x}_{0}$ is reduced. And then from this intermediate state we infer the desirable $\bm{x}_{0}$. There are several advantages of doing so: i) our solution is more efficient than the full reverse diffusion process from $\bm{x}_{T}$ to $\bm{x}_{0}$, ii) we do not need to retrain the diffusion model from scratch, and iii) we can still take advantage of the pre-trained diffusion model via the reverse Markov chain from $\bm{x}_{N}$ to $\bm{x}_{0}$.
SR3 [17] also exploits the potentials of the diffusion model for blind restoration. It feeds the LQ image into the diffusion model as a condition to guide the restoration at each timestep. This requires one to retrain the diffusion model from scratch on pre-collected training data. Hence, it would still suffer from the issue of degradation mismatch when dealing with real-world data. Unlike SR3, our method does not need to train the diffusion model from scratch but sufficiently leverages the prior knowledge of the pretrained diffusion model. The unique design on transition distribution $p(\bm{x}_{N}|\bm{y}_{0})$ further allows us to better cope with unknown degradations.
In summary, the contributions of this work are as follows:
  • We devise a new diffusion-based BFR approach to cope with severe and unknown degradations. Formulating the posterior distribution as a Markov chain that starts from the LQ image and ends at the desirable HQ image is novel. We show that the Markov chain can compress the predicted error by a factor of less than 1.
  • We show that the image prior captured in a pretrained diffusion model can be harnessed by having the Markov chain built partially on the reverse diffusion process. Such a unique design also allows us to explicitly control the restoration’s fidelity and realism by changing the Markov chain’s length.
  • We show that BFR can be achieved without complicated losses as in existing methods. Our method only needs to train a neural network with $L_{1}$ loss, simplifying the training pipeline. A preview of our results compared with existing methods is shown in Fig. 2.
II Related Work
The BFR approaches mainly focused on exploring better face priors. We review a few popular priors in this section, including geometric priors, reference priors, and generative priors. Diffusion prior can be considered as a type of generative prior.
Geometric Priors. Face images are highly structured compared with natural images. The structural information, such as facial landmarks [8], face parsing maps [18, 19, 18], facial component heatmaps [9], and 3D shapes [20, 21, 22], can be used a guidance to facilitate the restoration. The geometric face priors estimated from degraded inputs can be unreliable, affecting the performance of the subsequent BFR task.
Reference Priors. Some existing methods [23, 24] guide the restoration with an additional HQ reference image that owns the same identity as the degraded input. The main limitations of these methods stem from their dependence on the HQ reference images, which are inaccessible in some scenarios. Li et al. [9] address this limitation by constructing an offline facial component dictionary based on the features extracted from HQ images. Then, it searches for the closest facial components in this dictionary for the given LQ images during restoration.
Generative Priors. Unlike [9], more recent approaches directly exploit the rich priors encapsulated in generative models for BFR. Following the paradigm of GAN inversion [25], the earliest explorations [26, 27, 11, 28] iteratively optimize the latent code of a pretrained GAN for the desirable HQ target. To circumvent the time-consuming optimization, some studies [10, 1, 12] directly embed the decoder of the pre-trained StyleGAN [29] into the BFR network, and evidently improve the restoration performance. The success of VQGAN [30] in image generation also inspires several BFR methods. These methods mainly design different strategies, e.g., cross-attention [31], parallel decoder [4], and transformer [13], to improve the matching between the codebook elements of the degraded input and the underlying HQ image.
Attributed to the powerful generation capability of the diffusion model, some works based on diffusion model have been proposed recently. Typically, SR3 [17] and SRDiff [32] both feed the LQ image into the diffusion model as a condition to guide the restoration in training. To accelerate the inference speed, LDM [33] proposed to train the diffusion model in latent space of VQGAN [30]. These methods require one to retrain the diffusion model from scratch on some pre-collected data. Furthermore, the learned model would still be susceptible to the degradation mismatch when generalizing to other datasets. In a bid to circumvent the laborious and time-consuming retraining process, several investigations have explored the utilization of a pre-trained diffusion models as a generative prior to facilitate the restoration task, such as ILVR [34], RePaint [35], DDRM [36], CCDF [37], MCG [38], GDP [39] and so on [40, 41, 42, 43, 44]. The common idea underlying these approaches is to modify the reverse sampling process of the pre-trained diffusion model by introducing a well-defined or manually assumed degradation model as an additional constraint. Even though these methods perform well in certain ideal scenarios, they can not deal with the BFR task since its degradation model is unknown and complicated. To address these issues, the present study thus devises a new learning paradigm for BFR based on a pre-trained diffusion model.
III Preliminaries
Before presenting our method, we commence with a brief introduction to the diffusion probabilistic model, denoted as diffusion model for conciseness. Diffusion model can be traced back to its seminal exposition in [14], primarily inspired by the theory of non-equilibrium statistical physics. Ho et al. [16] reformulated this model from the perspective of denoising scoring matching, namely DDPM, achieving a huge success in the domain of image generation. In the course of this study, we follow the notations of DDPM for convenience.
Diffusion model consists of a forward process (or diffusion process) and a reverse process. Given a data point $\bm{x}_{0}$ with probability distribution $q(\bm{x}_{0})$, the forward process gradually destroys its data structure by repeated application of the following Markov diffusion kernel:
$q(\bm{x}_{t}|\bm{x_{t-1}})=\mathcal{N}(\bm{x}_{t};\sqrt{1-\beta_{t}}\bm{x}_{t-1},\beta_{t}\bm{I}),$ (1)
where $t\in\{1,2,\cdots,T\}$, $\{\beta_{t}\}_{t=1}^{T}$ is a pre-defined or learned noise variance schedule, and $\bm{I}$ is the identity matrix. With a rational design on the variance schedule, it theoretically guarantees that $q(\bm{x}_{T})$ converges to the unit spherical Gaussian distribution. It is noteworthy that the marginal distribution at arbitrary timestep $t$ has the following analytical form:
$q(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\sqrt{\alpha_{t}}\bm{x}_{0},(1-\alpha_{t})\bm{I}),$ (2)
where $\alpha_{t}=\prod_{l=1}^{t}(1-\beta_{l})$.
As for the reverse process, it aims to learn a transition kernel from $\bm{x}_{t}$ to $\bm{x}_{t-1}$, which is defined as the following Gaussian distribution:
$p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})=\mathcal{N}\left(\bm{x}_{t-1};\bm{\mu}_{\theta}(\bm{x}_{t},t),\sigma^{2}_{t}\bm{I}\right),$ (3)
where $\theta$ is the learnable parameter. With such a learned transition kernel, we can approximate the data distribution $q(\bm{x}_{0})$ via the following marginal distribution:
$p_{\theta}(\bm{x}_{0})=\int p(\bm{x}_{T})\prod_{t=1}^{T}p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})\mathrm{d}\bm{x}_{1:T},$ (4)
where $p(\bm{x}_{T})=\mathcal{N}(\bm{x}_{T};\bm{0},\bm{I})$.
In the framework of DDPM [16], the reparametrization of the reverse posterior $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$ is delineated as follows:
$\bm{\mu}_{\theta}(\bm{x}_{t},t)=\sqrt{\frac{\alpha_{t-1}}{\alpha_{t}}}\left(\bm{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\bm{\varepsilon}_{\theta}(\bm{x}_{t},t)\right),$
$\sigma^{2}_{t}=\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t},$ (5)
where $\bm{\varepsilon}_{\theta}(\bm{x}_{t},t)$ is designed to predict the noise contained in $\bm{x}_{t}$. Song et al. [45] generalize DDPM to a non-Markov process named DDIM, resulting in a deterministic generative process. Within this paradigm, the reverse posterior is expressed as:
$\bm{\mu}_{\theta}(\bm{x}_{t},t)=\sqrt{\alpha_{t-1}}\hat{\bm{x}}_{0}^{(t)}+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\bm{\varepsilon}_{\theta}(\bm{x}_{t},t),$
$\sigma^{2}_{t}=\eta\cdot\frac{1-\alpha_{t-1}}{1-\alpha_{t}}\beta_{t},~{}\eta\in[0,1],$ (6)
where
$\hat{\bm{x}}_{0}^{(t)}=\frac{\bm{x}_{t}-\sqrt{1-\alpha_{t}}\bm{\varepsilon}_{\theta}(\bm{x}_{t},t)}{\sqrt{\alpha_{t}}}.$ (7)
It is worth highlighting that the hyper-parameter $\eta$ controls the extent of randomness during inference. Specifically, when $\eta=0$, the reverse process becomes totally deterministic, while at $\eta=1$ it degenerates into the DDPM model.
Figure 3: Illustration of the diffused $\bm{x}_{N}$ (top row) and the reconstructed results (bottom row) by a pretrained diffusion model from different starting timesteps. Note that the employed diffusion model is trained with 1000 discrete steps following [15].
IV Proposed Method
In this section, we present the proposed BFR method DifFace that exploits the image priors encapsulated in a pretrained diffusion model. To keep the notations consistent with Sec. III, we denote the LQ image and HQ image as $\bm{y}_{0}$ and $\bm{x}_{0}$. To restore the HQ image from its degraded counterpart, we aim at designing a rational posterior distribution of $p(\bm{x}_{0}|\bm{y}_{0})$.
IV-A Motivation
Considering a diffusion model with $T$ discrete steps, it provides a transition function from $\bm{x}_{t}$ to $\bm{x}_{t-1}$. With the aid of this transition, we can construct the posterior distribution $p(\bm{x}_{0}|\bm{y}_{0})$ as follows:
$p(\bm{x}_{0}|\bm{y}_{0})=\int p(\bm{x}_{N}|\bm{y}_{0})\prod_{t=1}^{N}p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})\mathrm{d}\bm{x}_{1:N},$ (8)
where $1\leq N<T$ is an arbitrary timestep. Therefore, we can restore $\bm{x}_{0}$ from $\bm{y}_{0}$ by sampling from this posterior using ancestral sampling [46] from timestpe $N$ to $1$ as follows:
$\bm{x}_{N}\sim p(\bm{x}_{N}|\bm{y}_{0}),~{}~{}\bm{x}_{t-1}|\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t}).$ (9)
Since the transition kernel $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$ can be readily borrowed from a pretrained diffusion model, our goal thus turns to design the transition distribution of $p(\bm{x}_{N}|\bm{y}_{0})$.
We have an important observation by delving into Eq. (9). If replacing $p(\bm{x}_{N}|\bm{y}_{0})$ with the marginal distribution $q(\bm{x}_{N}|\bm{x}_{0})$ defined in Eq. (2), Eq. (9) degenerates into the diffusion and reconstruction process for $\bm{x}_{0}$ via a pretrained diffusion model, i.e.,
$\underbrace{\bm{x}_{N}\sim q(\bm{x}_{N}|\bm{x}_{0})}_{\text{Diffusion}},~{}~{}\underbrace{\bm{x}_{t-1}|\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})}_{\text{Reconstruction}}.$ (10)
In Figure 3, we show some diffused and reconstructed results under different settings for the starting timestep $N$. One can observe that when $N$ lies in a reasonable range (e.g., $N<500$), meaning that $\bm{x}_{0}$ is slightly “destroyed”, it is possible to accurately reconstruct $\bm{x}_{0}$ using the pretrained diffusion model.
This observation indicates that $q(\bm{x}_{N}|\bm{x}_{0})$ is an ideal choice for the desired $p(\bm{x}_{N}|\bm{y}_{0})$ by setting a reasonable $N$. Since the HQ image $\bm{x}_{0}$ is inaccessible in the task of BFR, we thus explore how to design a plausible $p(\bm{x}_{N}|\bm{y}_{0})$ to approximate $q(\bm{x}_{N}|\bm{x}_{0})$.
Figure 4: The curves of $\kappa_{N}$ and $\alpha_{N}$ with the starting timestep $N$.
IV-B Design
Recall that our goal is to design a transition distribution $p(\bm{x}_{N}|\bm{y}_{0})$ to approximate $q(\bm{x}_{N}|\bm{x}_{0})$. Fortunately, the target distribution $q(\bm{x}_{N}|\bm{x}_{0})$ has an analytical form as shown in Eq. (2). This inspires us to formulate $p(\bm{x}_{N}|\bm{y}_{0})$ as a Gaussian distribution as follows:
$p(\bm{x}_{N}|\bm{y}_{0})=\mathcal{N}\left(\bm{x}_{N};\sqrt{\alpha_{N}}f(\bm{y}_{0};w),(1-\alpha_{N})\bm{I}\right),$ (11)
where $f(\cdot;w)$ is a neural network with parameter $w$, aiming to provide an initial prediction for $\bm{x}_{0}$. It should be noted that the final restored result by our method is achieved by sampling from the whole Markov chain of Eq. (9) (see Fig. 1), and not directly predicted by $f(\cdot;w)$. As for $f(\cdot;w)$, it is only used to construct the marginal distribution of $\bm{x}_{N}$, a diffused version of $\bm{x}_{0}$, and thus named as “diffused estimator” in this work.
Next, we consider the Kullback-Leibler (KL) divergence between the designed $p(\bm{x}_{N}|\bm{y}_{0})$ and its target $q(\bm{x}_{N}|\bm{x}_{0})$. By denoting the predicted error of $f(\cdot;w)$ as $\bm{e}=\bm{x}_{0}-f(\bm{y}_{0};w)$, we have
$D_{\text{KL}}\left[p(\bm{x}_{N}|\bm{y}_{0})\|q(\bm{x}_{N}|\bm{x}_{0})\right]=\frac{1}{2}\kappa_{N}\|\bm{e}\|_{2}^{2},$ (12)
where $\kappa_{N}=\frac{\alpha_{N}}{1-\alpha_{N}}$. As shown in Fig. 4, $\kappa_{N}$ strictly decreases monotonically with the timestep $N$. Hence, larger $N$ will offer a better approximation to $q(\bm{x}_{N}|\bm{x}_{0})$, and further achieve a more realistic image via the designed posterior distribution of Eq. (8). However, $\bm{x}_{N}$ will contain more noises when $N$ is getting larger, as shown in Fig. 3. Thus, an overly large $N$ will inevitably deviate the restored result from the ground truth $\bm{x}_{0}$. Therefore, the choice of $N$ induces a realism-fidelity trade-off for the restored HQ image. We provide the ablation study in Sec. VI-B.
Algorithm 1 Inference procedure of DifFace
0:     The observed LQ image $\bm{y}_{0}$; The starting timestep $N$;
0:     The restored HQ image $\bm{x}_{0}$;
1:  Sample $\bm{x}_{N}$ from $p(\bm{x}_{N}|\bm{y}_{0})$ according to Eq. (11)
2:  for $t=N,N-1,\cdots,1$ do
3:     Refine $\hat{\bm{x}}_{0}^{(t)}$ via Eq. (16) (Optional)
4:     $\bm{x}_{t-1}=\sqrt{\alpha_{t-1}}\hat{\bm{x}}_{0}^{(t)}+\sqrt{1-\alpha_{t-1}}\bm{\varepsilon}_{\theta}(\bm{x}_{t},t)+\sigma_{t}\bm{\xi}$, where $\sigma_{t}$ is defined in Eq. (6), $\bm{\xi}\sim\mathcal{N}(\bm{\xi};\bm{0},\bm{I})$
5:  end for
IV-C Algorithm
Based on the constructed posterior distribution $p(\bm{x}_{0}|\bm{y}_{0})$ of Eq. (8), we can obtain the desired HQ image $\bm{x}_{0}$ through sequentially sampling from $p(\bm{x}_{N}|\bm{y}_{0})$ and $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$ for $t=N,\cdots,1$ as shown in Algorithm 1. However, it is essential to acknowledge that this stochastic sampling process inherently introduces potential deviation from the underlying HQ image, which is particularly problematic in the context of BFR that demands a high level of fidelity. To mitigate this fidelity issue, we utilize the DDIM sampler configured with $\eta=0.5$ when sampling from $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$. This strategy effectively eliminates the inherent randomness in the sampling process to some extent, enhancing the fidelity of the reconstructed HQ image. Detailed ablation studies on the settings of $\eta$ can be found in Sec. VI-B.
IV-D Discussion
Instead of directly learning a mapping from $\bm{y}_{0}$ to $\bm{x}_{0}$ under multiple constraints like current deep learning-based methods [18, 12, 1], the proposed method circumvents this by predicting $\bm{x}_{N}$, a diffused version of $\bm{x}_{0}$, from $\bm{y}_{0}$. Such a new learning paradigm brings several significant advantages compared to existing approaches:
  • Error Contraction. Considering the diffused estimator $f(\cdot;w)$, its predicted error is denoted as $\bm{e}=\bm{x}_{0}-f(\bm{y}_{0};w)$. In our method, the problem of BFR is reformulated to predict $\bm{x}_{N}$ as explained in Sec. IV-A. According to Eq. (11), $\bm{x}_{N}$ can be accessed as follows:
    $\bm{x}_{N}$ $=\sqrt{\alpha_{N}}f(\bm{y}_{0};w)+\sqrt{(1-\alpha_{N})}\bm{\zeta}$
    $=\sqrt{\alpha_{N}}\bm{x}_{0}-\sqrt{\alpha_{N}}\bm{e}+\sqrt{(1-\alpha_{N})}\bm{\zeta},$ (13)
    where $\bm{\zeta}\sim\mathcal{N}(\bm{\zeta}|\bm{0},\bm{I})$. It can be seen that the predicted error $\bm{e}$ is contracted by a factor of $\sqrt{\alpha_{N}}$ under our proposed paradigm of predicting $\bm{x}_{N}$, where $\sqrt{\alpha_{N}}$ is less than 1 as shown in Fig. 4.
    Attributed to such an error contraction, our method has a greater error tolerance to the diffused estimator. Hence, $f(\cdot;w)$ can be simply trained with $L_{1}$ loss on some synthetic data in this work. This unique flexibility bypasses the sophisticated training process compared with most recent methods [9, 1, 4]. What is more, this capability of compressing errors intuitively improves the robustness of our method, especially when dealing with severe and complex degradations.
  • Diffusion Prior. After obtaining the diffused $\bm{x}_{N}$ via Eq. (13), our method gradually generates the desirable HQ result by sampling recursively from $p_{\theta}(\bm{x}_{t-1}|\bm{x}_{t})$ starting from $t=N$ and ending at $t=1$. Through this sampling procedure, we effectively leverage the rich image priors and powerful generation capability of the pretrained diffusion model to help the restoration task. Unlike existing methods, since the diffusion model is completely trained on the HQ images in an unsupervised manner, it thus reduces the dependence of our method on the manually synthesized training data, of which the distribution may deviate from the true degradation.
These two intrinsic properties are mainly delivered by the constructed posterior in Eq. (8), which is the core formulation of our method. Such a posterior is specifically designed for BFR to mitigate the robustness issue caused by complicated degradations. Therefore, our proposed method indeed renders a new simple yet robust learning paradigm for BFR, rather than a straightforward application of diffusion model in this problem.
V Extension
The proposed DifFace constitutes a versatile framework, being able to easily extended to a variety of blind image restoration tasks, such as super-resolution, deblurring and so on. The extensibility of DifFace can be readily achieved through a straightforward substitution of the training data to train the diffused estimator. This section delves into a distinct facet of non-blind restoration, typified by a well-defined degradation model, e.g., inpainting and colorization. The availability of the degradation model in these scenarios enables us to incorporate an additional optimization strategy to further improve the efficacy of our method. For clarity and pedagogical purposes, we take the task of inpainting as an example to elucidate the precise technical details involved.
The goal of inpainting is to restore the underlying HQ image $\bm{x}_{0}$ given the observed LQ image $\bm{y}_{0}$, while adhering to the following constraint, i.e.,
$\bm{y}_{0_{[i,j]}}=\begin{dcases}\bm{x}_{0_{[i,j]}},&\text{if}~{}\bm{M}_{[i,j]}=0,\\ 0,&\text{if}~{}\bm{M}_{[i,j]}=1,\end{dcases}$ (14)
where $\bm{A}_{[i,j]}$ denotes the element located at position $(i,j)$ of the matrix $\bm{A}$, and $\bm{M}$ represents the binary image mask filled with 00 or $1$. In the inference procedure of DifFace, we generate $\bm{x}_{t-1}$ from $\bm{x}_{t}$ based on the estimated value for $\bm{x}_{0}$, namely $\hat{\bm{x}}_{0}^{(t)}$, in each timestep as exemplified in Eq. (6). This inspires us to refine $\hat{\bm{x}}_{0}^{(t)}$ according to the degradation constraint of Eq. (14). By denoting the refined target as $\ddot{\bm{x}}_{0}^{(t)}$, the refinement process can be formulated as an optimization problem:
$\min_{\ddot{\bm{x}}_{0}^{(t)}}\left\|\ddot{\bm{x}}_{0}^{(t)}-\hat{\bm{x}}_{0}^{(t)}\right\|_{2}+\gamma\left\|(\bm{1}-\bm{M})\odot\left(\bm{y}_{0}-\ddot{\bm{x}}_{0}^{(t)}\right)\right\|_{2},$ (15)
where $\odot$ denotes the Hadamard product, $\gamma$ is a hyper-parameter controlling the balance between the diffusion prior and the degradation constraint. Furthermore, this problem admits a closed-form solution as follows [^1] All the mathematical operations in Eq (16) are pixel-wise.:
$\ddot{\bm{x}}_{0}^{(t)}=\frac{\hat{\bm{x}}_{0}^{(t)}+\gamma(\bm{1}-\bm{M})^{2}\odot\bm{y}_{0}}{\bm{1}+\gamma(\bm{1}-\bm{M})^{2}}.$ (16)
It should be noted that in cases where a closed-form solution is not feasible, Eq. (16) can be substituted with a single stochastic gradient descent (SGD) step.
Figure 5: An example restored by DifFace under different settings of the starting timestep $N$ and the diffused estimator $f(\cdot;w)$. (a) LQ image, (b) HQ image, (c1)-(c2) restored results by SRCNN and SwinIR, (d1)-(i1) restored results by DifFace that takes SRCNN as a diffused estimator, (d2)-(i2) restored results by DifFace that takes SwinIR as a diffused estimator.
[…]
Methods
Metrics
9
[…]
# Parameters(M)
SRCNN
23.58
0.722
0.555
73.09
11.13
133.49
108.68
1.03
[…]
23.03
0.670
0.517
70.80
8.40
58.53
44.83
[…]
SwinIR
24.32
0.736
0.507
63.80
5.82
131.82
93.73
15.79
[…]
23.44
0.670
0.461
64.94
6.06
48.98
20.29
[…]
Figure 6: From left to right: (a) FID and LPIPS with respect to the starting timestep $N$ on the validation dataset, (b) LPIPS with respect to the downsampling factor, namely $s$ in Eq. (17), on the testing dataset.